Question 6. Consider a perfectly competitive market. The equilibrium price is 12. The equilib- rium quantity is 440. The supply curve is Q (P ) = 200 + 20P . Which of the following could be the demand curve in this market?
a) Noneoftheotheranswersiscorrect.
b) Q(P)=500−5P
c) Q(P)=300−10P
d) Q(P)=1000−10P
1
Expert's answer
2019-04-10T09:10:42-0400

Let's calculate the quantity corresponding to each of the equations b), c), d) at the equilibrium price of 12. So, we get:

b) Q(P)=500−5P=500-5*12=440,

c) Q(P)=300−10P=300-10*12=180,

d) Q(P)=1000−10P=1000-10*12=880.

The equilibrium quantity is 440 that corresponds to the option b).

Answer: The option b) could be the demand curve equation in this market.